### v2

# Avoiding Catastrophe in Continuous Spaces by Asking for Help

Abstract
Most reinforcement learning algorithms with formal regret guarantees assume all mistakes are reversible and essentially rely on trying all possible behaviors. This approach leads to poor outcomes when some mistakes are irreparable or even catastrophic. We propose a variant of the contextual bandit problem where the goal is to minimize the chance of catastrophe. Specifically, we assume that the payoff each round represents the chance of avoiding catastrophe that round, and try to maximize the product of payoffs (the overall chance of avoiding catastrophe). We allow a limited number of queries to a mentor and assume a Lipschitz continuous payoff function. We first show that in general, any algorithm either constantly queries the mentor or is nearly guaranteed to cause catastrophe. However, when the mentor policy class has bounded Natarajan dimension and contains at least some “reasonable” policies, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows. We also present an alternative algorithm which provides the same regret and query guarantees when the mentor’s action changes a constant number of times in a 1D state space, and can handle adversarially chosen states.
1 Introduction
In reinforcement learning (RL), an agent finds itself in an unknown environment and wants to maximize some payoff function, commonly modeled as a Markov Decision Process (MDP). Typically, the only way for the agent to learn about its environment is by trying actions and observing their effects. This presents a challenge when some actions are inherently dangerous: for example, the agent must learn to avoid self-destructive actions without actually trying those actions.
This problem is hopeless without additional assumptions, since any untried action could lead to either paradise or disaster, and the agent has no way to predict which. Most RL theory evades this issue by either assuming ergodicity, meaning that any state is reachable from any other (e.g., [15]), or by resetting the agent at the start of each “episode” (e.g., [1]). In the real world, however, many actions lead to unrecoverable situations: we call these catastrophes. Examples could include ingesting poison or crashing a robotic vehicle. Resetting the agent may be expensive (if it destroys itself), inadequate (if it irreparably harms others), or impossible (if it is outside of human supervision).
Rather than ignore the possibility of catastrophe, we explore whether access to a mentor might enable an agent to navigate risky (i.e., non-ergodic single-episode) environments. We assume the mentor illustrates the optimal action in the current state, and we would like an algorithm which obtains near-optimal payoff with a limited number of mentor queries. “Near-optimal payoff” is often modeled via regret, defined as the gap between the optimal total payoff and the agent’s actual total payoff. For undiscounted MDPs (i.e., each time step has the same weight in the objective), one typically aims for regret sublinear in the time horizon $T$, or equivalently, for the average regret per time step to approach 0 for large $T$. Similarly, a sublinear number of mentor queries implies that the rate of querying the mentor goes to 0 for large $T$—that is, the agent gradually becomes self-sufficient. Formally, we pose the following open problem:
Is there an algorithm for non-ergodic single-episode undiscounted MDPs which ensures that both the regret and the number of mentor queries are sublinear in $T$?
An additional assumption is necessary, as the agent needs some way to make inferences about unseen states in order to decide when to query. Much past work has used Bayesian inference, which suffers tractability issues in complex environments. [^1] For the curious reader, [3] provides a thorough treatment. See also Section 1.3 for further discussion. In this paper, we use Lipschitz continuity instead, which captures the intuition that an action should have similar effects in similar states. Importantly, the Lipschitz assumption is compatible with any state space and action space with a distance metric.
1.1 Our model
In this paper, we target a more modest goal: can the agent simply avoid catastrophe? This is clearly necessary (but not sufficient) for sublinear regret in general MDPs. Furthermore, avoiding catastrophe is the precise ability which standard RL algorithms lack.
Focusing on catastrophe also allows us to streamline the model. If we redefine the payoff at time $t$ to be the probability of avoiding catastrophe on that time step, then the product of payoffs represents the overall chance of avoiding catastrophe. Representing catastrophe in the objective function eliminates the need to model non-ergodicity through the transition function, as MDPs do. Our multiplicative objective captures the same stringency: a single mistake can ruin the entire endeavor. [^2] An additive objective with unbounded payoffs could also capture the core challenge of needing to avoid catastrophic actions without ever trying them. However, we feel that the multiplicative objective is more intuitive due its interpretation as the probability of avoiding catastrophe.
The resulting model is similar to the contextual bandit problem: on each time step, the agent observes a state (we generally assume i.i.d. states), selects an action (or queries the mentor), and obtains some payoff. To our knowledge, multiplicative objective functions have not been studied for contextual bandits. Lastly, unlike the standard contextual bandit model, we assume that the agent does not observe payoffs. This is because the payoff in our model represents the chance of avoiding catastrophe that round, and in the real world, one only observes whether catastrophe occurred, not its probability. [^3] One may be able to detect “close calls” in some cases, but observing the precise probability seems unrealistic.
1.2 Our results
We first show that in general, any algorithm with sublinear queries to the mentor has arbitrarily poor regret in the worst-case (Theorem 3.1). As a corollary, even when it is possible to avoid catastrophe with certainty, any algorithm either asks for help on a constant fraction of time steps or is nearly guaranteed to cause catastrophe. This holds even for randomized algorithms and simple environments: our construction uses just two actions and assumes states are uniformly random across $[0,1]$.
On the other hand, the class of mentor policies in our negative result is quite complex in the sense of having unbounded VC dimension [31]. What if we restrict the mentor policy to be relatively simple?
We present an intuitive algorithm whose regret and rate of querying the mentor both go to 0 as $T\to\infty$ when the mentor policy class has bounded VC dimension (or Natarajan dimension, more generally [^4] Natarajan dimension [24] generalizes VC dimension (which only handles binary prediction) to multi-class prediction. In our case, each class is an action, and a policy is a multi-class predictor from states to actions.) and contains at least some “reasonable” policies (Theorem 4.5). Our algorithm can handle an arbitrary unknown distribution over states, a multi-dimensional unbounded state space, and many actions. It does not need detailed access to the feature embedding (instead relying on two simple operations) and does not need to know the Lipschitz constant. It does need to know the policy class and (an upper bound on) the Natarajan dimension (as is standard in PAC learning). This is our primary positive result.
Finally, we give another algorithm with the same guarantees of sublinear queries and near-zero regret under a different simplicity assumption (Theorem 5.2). Specifically, we assume that the mentor’s action changes a constant number of times in a 1D state space. This algorithm can also handle adversarially chosen states.
We generally assume that the mentor’s policy is optimal, but our positive results can be generalized to a suboptimal mentor under a slightly different Lipschitz assumption (Remark B.1).
1.3 Related work
See [9; 11] for full surveys of safe RL. Here, we discuss only the work most relevant to our own.
Learning with irreversible costs.
Despite the ubiquity of irreversible costs in the real world, there remains a dearth of theoretical work on this topic. This may be due to the fundamental modeling question of how the agent should learn about novel states or actions without actually trying them.
The most common approach is to allow the agent to ask for help. This alone is insufficient, however: the agent must have some way to decide when to ask for help. A popular solution is to perform Bayesian inference on the world model, but this has two tricky requirements: (1) a prior distribution which contains the true world model (or an approximation), and (2) an environment where computing (or approximating) the posterior is tractable. A finite set of possible environments satisfies both conditions, but is unrealistic in many real-world scenarios. In contrast, our algorithm can handle a continuous unbounded state space, which is crucial for many real-world scenarios in which one never sees the exact same state twice.
Bayesian inference combined with asking for help is studied by [5; 6; 17; 21]. We also mention [12; 22; 30], which utilize Bayesian inference in the context of safe RL without asking for help (and without regret bounds).
We are only aware of two papers which theoretically address irreversibility without Bayesian inference: [10] and [20]. The former proposes to sample trajectories and learn reversibility based on temporal consistency between states: intuitively, if $s_{1}$ always precedes $s_{2}$, we can infer that $s_{1}$ is unreachable from $s_{2}$. Although the paper theoretically grounds this intuition, there is no formal regret guarantee. The latter presents an algorithm which asks for help in the form of rollouts from the current state. However, the regret bound and number of rollouts are both linear in the worst case, due to the dependence on the $\gamma^{*}$ parameter which roughly captures how bad an irreversible action can be. In contrast, our algorithms achieve near-zero regret even when actions can be maximally bad.
To our knowledge, we are the first to provide an algorithm which formally guarantees avoidance of catastrophe (with high probability) without Bayesian inference.
Contextual bandits.
Like MDPs, the contextual bandit model involves an agent who selects an action on each time step based on the current state (or “context”) to maximize payoff. Unlike MDPs, the state on each time step is determined randomly or by an adversary, instead of by a transition function. See [29] for an overview.
The contextual bandit literature often uses Lipschitz continuity of payoff functions as a way to generalize between states, which affirms our use of that same assumption [4; 14; 18; 19; 26; 28]. However, to our knowledge, all prior contextual bandit work aims to maximize the sum of payoffs, which implicitly assumes that no mistake is catastrophic: no matter how poorly you perform, you can always make it up given a sufficiently long time horizon.
Multiplicative objectives.
Perhaps the best-studied example of a multiplicative objective is the Nash welfare function (also known as the Nash bargaining solution), defined as the geometric mean of agent utilities [16; 23]. Inspired by Nash welfare, [2] recently studied algorithms for maximizing the geometric mean of payoffs across time steps in the classic (non-contextual) multi-armed bandit problem. However, the geometric mean of payoffs is much more forgiving of errors compared to the product of payoffs. Imagine that the agent obtains a payoff of $1-1/\sqrt{T}$ per time step: then $\lim_{T\to\infty}(1-1/\sqrt{T})^{T}=0$ but $\lim_{T\to\infty}((1-1/\sqrt{T})^{T})^{1/T}=\lim_{T\to\infty}(1-1/\sqrt{T})=1$. In our setting, it is crucial that the cumulative chance of catastrophe (the product of payoffs) is small, not just the average chance of catastrophe per time step (the geometric mean of payoffs).
Multiplicative objectives also arise in the context of compound growth or repeated gambles, where one’s wealth at time $t$ can be modeled as the initial wealth times a product of growth factors (see, e.g., [25]). Our problem could be viewed as a compound growth problem where one’s “wealth” (the total probability of avoiding catastrophe) never increases. However, we are not aware of any mathematical tools related to compound growth that handle the contextual bandit aspects of our problem.
The rest of the paper is structured as follows. Section 2 formally defines our model. Section 3 presents our negative result for general mentor policies. Section 4 presents our positive result for policy classes with bounded Natarajan dimension. Section 5 presents our second positive result under a different simplicity assumption. Section 6 discusses limitations. All proofs are deferred to the appendix.
2 Model
Lipschitz contextual bandits.
Let $\mathbb{N}$ refer to the strictly positive integers and let $T\in\mathbb{N}$ be the time horizon. Let $S\subseteq\mathbb{R}^{n}$ be the state space, and let $\mathbf{s}=(s_{1},s_{2},\dots s_{T})\in S^{T}$ be a sequence of states. (The contextual bandit literature typically uses “context” instead of “state”, but we have chosen to use “state” to align with MDP terminology.) With the exception of Section 5, we assume that each state is sampled i.i.d. from an unknown distribution $\mathcal{D}$ over $S$.
For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$. On each time step $t\in[T]$, the agent must select an action $a_{t}\in A\cup\{\hat{a}\}$, which generates payoff $\mu(s_{t},a_{t})\in[0,1]$. We allow the agent’s choice of action to be randomized. The action $\hat{a}$ constitutes querying the mentor, which we will define shortly. The agent never observes payoffs: the only way to learn about $\mu$ is by querying the mentor. We assume that $|A|\leq T$, since the agent can never observe more than $T$ distinct actions.
We assume that $\mu$ is Lipschitz continuous with respect to $S$: there exists $L>0$ such $\forall s,s^{\prime}\in S,a\in A$, $|\mu(s,a)-\mu(s^{\prime},a)|\leq L||s-s^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance.
Asking for help.
The special action $\hat{a}$ constitutes querying a mentor who knows an optimal policy $\pi^{*}:S\to A$. When action $\hat{a}$ is chosen, the mentor informs the agent of the action $\pi^{*}(s_{t})$ and the agent obtains optimal payoff $\mu(s_{t},\pi^{*}(s_{t}))=\mu^{*}(s_{t})=\max_{a\in A}\mu(s_{t},a)$. Let $Q_{t}=\{i\in[t]:a_{i}=\hat{a}\}$ be the set of time steps up to and including $t$ which involve a query to the mentor.
We would like an algorithm which becomes “self-sufficient” over time: the rate of querying the mentor should go to 0 as $T\to\infty$. Intuitively, this is equivalent to the cumulative number of queries being sublinear in $T$: $|Q_{T}|\in o(T)$. However, interpreted literally, $|Q_{T}|\in o(T)$ is not a rigorous definition. This is because $|Q_{T}|$ is not just a function of $T$: it can also depend on $\mathbf{s}$, $\mu$ and $\pi^{*}$.
To formalize this, let $|Q_{T}(\mathbf{s},\mu,\pi^{*})|$ be the number of queries to the mentor up to time $T$, given payoff function $\mu$ and state sequence $\mathbf{s}$. Then we say that $|Q_{T}|$ is sublinear in $T$, denoted $|Q_{T}|\in o(T)$, if there exists $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $|Q_{T}(\mathbf{s},\mu,\pi^{*})|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{*}$.
Multiplicative objective and regret.
If $\mu(s_{t},a_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu(s_{t},a_{t})$ represents the agent’s overall chance of avoiding catastrophe. [^5] Conditioning on no prior catastrophe means we do not need to assume that these probabilities are independent (and if catastrophe has already occurred, this time step does not matter). This is due to the chain rule of probability. For a time horizon $T$, payoff function $\mu$, mentor policy $\pi^{*}$, and state sequence $\mathbf{s}$, the agent’s regret is
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
where $\mathbf{a}=\{a_{1},\dots,a_{T}\}$ is the set of actions selected by the algorithm (which depends on $\pi^{*}$). We focus on a worst-case choice of $\mu$ and $\pi^{*}$. [^6] The choice of $\pi^{*}$ can actually matter: suppose $a$ is optimal everywhere but $a^{\prime}$ is optimal only for this state.
VC and Natarajan dimension.
VC dimension [31] and Natarajan dimension [24] are standard measures of learning difficulty. The former applies to binary classes, while the latter handles multi-class prediction (and reduces to the former in the case of two classes). In our setting, a policy can be viewed as a multi-class predictor, where each action is a class. The precise definitions of these dimensions are not relevant to our paper, as we only invoke Natarajan dimension indirectly through the fundamental theorem of multi-class PAC learning (Theorem 4.4). We direct the reader to [27] for a comprehensive overview.
Misc.
We will often deal with the measure of subsets of $S$. When $n=1$, we use $\operatorname*{len}(X)$ to denote the length of interval $X\subset S$; for general $n$, we use $\operatorname*{vol}(X)$ to denote the $n$-dimensional measure of $X\subseteq S$ [^7] Although we could use the same notation for both cases, we find this to be more intuitive. The convex hull of a set $X$ is denoted $\operatorname*{conv}(X)$. All logarithms and exponents are base $e$ unless otherwise noted.
3 Hardness result for complex mentor policies
We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when states are uniformly random on $[0,1]$ and there are only two actions. The result also holds even if the algorithm knows $L$ and $\mathbf{s}$ ahead of time.
Let $U(S)$ denote the uniform distribution on $S$ and let $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\mathbf{a}}$ denote the expectation over both $\mathbf{s}$ and any additional randomness in the algorithm.
Theorem 3.1.
The worst-case expected regret of any algorithm with sublinear queries goes to 1 as $T$ goes to infinity. Formally,
$\lim_{T\to\infty}\ \,\sup_{\mu,\pi^{*}}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ \,[R_{T}(\mathbf{s},\mu,\pi^{*})]=1$
3.1 Intuition
We partition $S$ into equally-sized sections that are “independent” in the sense that querying a state in section $i$ gives you no information about section $j$. The number of sections is determined by a function $f:\mathbb{N}\to\mathbb{N}$ that we will choose. If $|Q_{T}|\in o(f(T))$, most of these sections will never contain a query. When the agent sees a state in a section not containing a query, it essentially has to guess, meaning it will be wrong a constant fraction of the time. Figure 1 fleshes out this idea.
Picking $\boldsymbol{f(T)}$.
A natural idea is to try $f(T)=T$. Unfortunately, this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$. In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$. Lemma A.1 implies the existence of such an $f(T)$ if $|Q_{T}|$ is bounded by some function $g(T)\in o(T)$, so we get a bound of the following form: $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$. [^8] The bound in the actual proof will be in log space, so it will look slightly different, but the idea is the same. When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})=1$. In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
$s$$\mu$  $\mu(s,0)$$\mu(s,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
Figure 1: An illustration of the construction we use to prove Theorem 3.1 (not to scale). The horizontal axis indicates the state $s\in[0,1]$ and the vertical axis indicates the payoff $\mu(s,a)\in[0,1]$. The solid line represents $\mu(s,0)$ and the dotted line represents $\mu(s,1)$. In each section, one of the actions has the optimal payoff of 1, and the other action has the worst possible payoff allowed by the Lipschitz constraint, reaching a minimum of $1-\frac{L}{2f(T)}$. Crucially, both actions result in a payoff of 1 at the boundaries between sections: this allows us to “reset” for the next section. As a result, we can freely toggle the optimal action for each section independently.
VC/Natarajan dimension.
The class of mentor policies induced by our construction has VC and Natarajan dimensions of $f(T)$. [^9] Recall that VC and Natarajan dimensions coincide for $|A|=2$. This is expected given our positive result for policy classes with bounded Natarajan dimension (Theorem 4.5).
3.2 Formal definition of construction
Let $S=[0,1]$, $A=\{0,1\}$, and $\mathcal{D}=U(S)$. Assume that $L\leq 1$; this will simplify the math and only makes the problem easier for the agent. We define a family of payoff functions parameterized by a function $f:\mathbb{N}\to\mathbb{N}$ and a bit string $\mathbf{x}=(x_{1},x_{2},\dots,x_{f(T)})\in\{0,1\}^{f(T)}$. The bit $x_{j}$ will denote the optimal action in section $j$. Note that $f(T)\geq 1$ and since we defined $\mathbb{N}$ to exclude 0.
[…]
$\mu_{f,\mathbf{x}}(s,a)=\begin{cases}1&\textnormal{ if }a=x_{j(s)}\\ 1-L\left(\cfrac{1}{2f(T)}-|m_{j(s)}-s|\right)&\textnormal{ if }a\neq x_{j(s)}\\ \end{cases}$
For any $\mathbf{x}\in\{0,1\}^{f(T)}$, $\mu_{f,\mathbf{x}}$ is piecewise linear (trivially) and continuous (because both actions have payoff 1 on the boundary between sections). Since the slope of each piece is in $\{-L,0,L\}$, $\mu_{f,\mathbf{x}}$ is Lipschitz continuous with Lipschitz constant $L$.
Since each $s_{t}$ belongs to exactly one $S_{j}$, there is a unique optimal action for each $s_{t}$, so the choice of $\pi^{*}$ does not matter in this scenario.
4 Near-zero regret assuming bounded Natarajan dimension
In this section, we present a simple algorithm with near-zero regret when the policy class $\Pi$ has bounded Natarajan dimension $d$ and which contains at least some “reasonable” policies. The algorithm needs to know $\Pi$ and (an upper bound on) $d$, as is standard in PAC learning (see, e.g., [27]). The algorithm does not need to know $L$, and can handle an unbounded state space (the number of queries simply scales with the volume of the convex hull of the observed states).
We need a few definitions to formally state our algorithm and result.
Definition 4.1.
A policy $\pi$ is consistent with $X\subset S\times A$ if $\pi(s)=a$ for all $(s,a)\in X$.
Definition 4.2.
Given $X\subset S\times A$ and $r\geq 0$, a policy $\pi$ is $(X,r)$-local if (1) $\pi$ is consistent with $X$ and (2) whenever $(s^{\prime},a^{\prime})\in X$ and $||s-s^{\prime}||\leq r$, there exists $(s^{\prime\prime},a^{\prime\prime})\in X$ such that $||s-s^{\prime\prime}||\leq r$ and $\pi(s)=a^{\prime\prime}$.
A policy which knows $X$ but is not $(X,r)$-local says either (1) “even though I know the optimal action for $s$, I’m going to do something else” or (2) “even though I have a data point near $s$, I’m going to base my action on a data point far from $s$”. [^10] Note that the mentor policy cannot itself query the mentor, so we are only dealing with actions in $A$. Given the Lipschitz nature of the problem, this type of policy seems quite unintuitive. However, we allow $\Pi$ to contain non-local policies, as long as it also contains local policies. Note that knowing $X,r$, and $\Pi$ is sufficient to find a $(X,r)$-local policy.
Definition 4.3.
Given $r\geq 0$, $\Pi$ is $r$-local if for every $\tilde{S}\subset S$ and $X=\{(s,\pi^{*}(s)):s\in\tilde{S}\}$, there exists an $(X,r)$-local policy in $\Pi$.
As an example, let $S=\mathbb{R}$ and suppose $\Pi$ is the set of policies where $\pi(s)$ changes at most $K$ times as $s$ goes from $\infty$- ∞ to +$\infty$+ ∞ (in the language of Section 5, $\pi$ has at most $K$ segments). Then $\Pi$ is $r$-local for any $r$, since for any set of queries $X$, the nearest neighbor policy is $(X,r)$-local and has at most $K$ segments (and thus is in $\Pi$). Note that this $\Pi$ also has bounded Natarajan dimension.
4.1 Intuition behind the algorithm
Our algorithm has a simple structure: gather enough data to choose a solid baseline policy, and then mostly follow that policy, but ask for help if you find yourself out-of-distribution (OOD). The algorithm does not need direct access to the feature embedding of states: the only two operations it uses are (1) update baseline policy and (2) find nearest neighbor distance. This approach seems quite natural for real-world applications, where one can imagine updating the baseline policy via gradient descent and treating an OOD detector score as the nearest neighbor distance.
In our case, the initial dataset $X$ is gathered by querying on the first $g(T)$ time steps. We will choose $g(T)$ to be large enough so that – along with the assumption on Natarajan dimension – we can apply the fundamental theorem of multi-class PAC learning (Theorem 4.4). [^11] Theorem 4.4 is called the “realizable” case of PAC learning, in contrast to the “agnostic” case where $\pi^{*}$ may not be in $\Pi$. Some sources (e.g., [27]) state that $\pi$ should minimize empirical loss, rather than being consistent with $X$. These are equivalent in the realizable case since consistent policies have zero empirical loss. This is the only place in the proof where Natarajan dimension is used.
Theorem 4.4 (Fundamental Theorem of Multi-Class PAC Learning [13; 24; 27]).
Assume $\Pi$ has finite Natarajan dimension $d$ and $\pi^{*}\in\Pi$. Then $\exists\,c>0$ such that $\forall\,\varepsilon,\delta\in(0,1)$, if $X\subset S\times A$ consists of at least $\frac{c}{\varepsilon}\left(d\log(\frac{|A|d}{\varepsilon})+\log\frac{1}{\delta}\right)$ i.i.d. samples $\mathcal{D}$ and the corresponding actions from $\pi^{*}$, then with probability at least $1-\delta$, any policy $\pi\in\Pi$ consistent with $X$ satisfies $\Pr_{s\sim\mathcal{D}}[\pi(s)\neq\pi^{*}(s)]\leq\varepsilon$.
Theorem 4.4 will ensure that $\Pr[a_{t}\neq\pi^{*}(s_{t})]$ is small, but even a single incorrect action can spell defeat if the payoff is bad enough. To handle this, the algorithm continues to selectively ask for help, but only when the state is far from any previously queried state (i.e., out-of-distribution). The policy $\pi$ is then updated based on this new information. The threshold $r$ is chosen such that we can cover the entire state $S$ with a sublinear number of queries.
1:function LookBeforeYouLeap($T\in\mathbb{N},\>r\in\mathbb{R}_{\geq 0},\>g:\mathbb{N}\to\mathbb{N}$)
2:     $X\leftarrow\emptyset$
3:     for $t$ from $1$ to $g(T)$ do $\triangleright$ Gather baseline data set
4:         Query mentor and observe $\pi^{*}(s_{t})$
5:         $X\leftarrow X\cup\{(s_{t},\pi^{*}(s_{t}))\}$
6:     $\pi\leftarrow$ any $(X,r)$-local policy in $\Pi$
7:     for $t$ from $g(T)+1$ to $T$ do $\triangleright$ Mostly follow $\pi$, but ask for help if out-of-distribution
8:         if $\min_{(s,a)\in X}||s_{t}-s||\leq r$ then
9:              Take action $\pi(s_{t})$
10:         else
11:              Query mentor and observe $\pi^{*}(s_{t})$
12:              $X\leftarrow X\cup\{(s_{t},\pi^{*}(s_{t}))\}$
13:              $\pi\leftarrow$ any $(X,r)$-local policy in $\Pi$
Algorithm 1 achieves near-zero regret when $\Pi$ has bounded Natarajan dimension and is $r$-local.
Tractability. Our algorithm can only be useful in practice if finding a $(X,r)$-local policy in $\Pi$ is computationally tractable. Such guarantees are impossible without further assumptions on $\Pi$ (this is again standard for PAC learning). However, if $\Pi$ supports nearest-neighbor policies (which satisfy locality for all values for $r$), those policies are both intuitive and can be computed easily.
We are now ready to state the main result of this section. The constant $c$ in Theorem 4.5 is the same as the constant $c$ in Theorem 4.4.
Theorem 4.5.
[…]
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq\frac{6L}{T^{\frac{0.5}{n+0.5}}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
Corollary 4.6 makes the near-zero regret and sublinear mentor queries more explicit.
Corollary 4.6.
Assume $\pi^{*}\in\Pi$, where $\Pi$ has finite Natarajan dimension $d$ and is $\varepsilon(T)^{1/n}$-local for $\varepsilon(T)=T^{\frac{-n}{n+0.5}}$. Then Algorithm 1 with $g(T)=2cdT^{\frac{n}{n+0.5}}\log(d\,T)$ and $r=\varepsilon(T)^{1/n}$ makes at most $O\left(T^{\frac{n}{n+0.5}}(d\log(d\,T)+\operatorname*{vol}(\operatorname*{conv}(S)))\right)$ queries and satisfies
$\lim_{T\to\infty}\>\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]=0$
When $d$ is constant, the number of queries in Corollary 4.6 is $o(T)$. (Actually, $d$ can be as large as $T^{\frac{0.5-\delta}{n+0.5}}$ for $\delta\in(0,1/2)$.) Also, since the expected regret goes to zero, the regret must go to zero with high probability (since regret is always nonnegative).
Finally, note that standard PAC learning lower bounds do not directly yield lower bounds for our setting: even $\Pr[a_{t}\neq\pi^{*}(s_{t})]=1$ is insufficient for high regret, since the payoffs could be nearly optimal. Also, PAC learning lower bounds typically state that there exists a distribution for which the algorithm performs poorly, while our lower bound uses a fixed known distribution.
5 Near-zero regret for simple mentor policies in 1D with adversarial states
Our final result is another algorithm with near-zero regret and sublinear queries, but under different assumptions. In this section, we assume a 1D state space and we allow the state sequence to be adversarially chosen. We consider on the following notion of simplicity:
Definition 5.1.
Given a mentor policy $\pi^{*}$, partition the state space $S$ into intervals such that all states within each interval share the same mentor action. Let $\{S_{1},\dots,S_{k}\}$ be a partition that minimizes the number of intervals. We call each $S_{j}$ a segment. Let $f(\pi^{*})$ denote the number of segments in $\pi^{*}$.
This is similar conceptually to Natarajan dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., optimal actions) on $\mathbf{s}$. The number of segments also relates to the minimum depth needed by a neural network to learn the policy [7]. Figure 1 shows why the Lipschitz constant does not bound the number of segments ($f(T)$ can be arbtrarily large).
We prove the following result:
Theorem 5.2.
For any $c\in[0,1]$, any $\mathbf{s}\in S^{T}$, and any $\pi^{*}$ with $f(\pi^{*})\leq K$, Algorithm 2 with $g(T)=T^{c}$ makes at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries and satisfies $R_{T}(\mathbf{s},\mu,\pi^{*})\leq 6LKT^{1-2c}$.
Corollary 5.3 makes the sublinear queries and near-zero regret of Theorem 5.2 more explicit when the number of segments is a constant.
Corollary 5.3.
For any $c\in(1/2,1)$, Algorithm 2 with $g(T)=T^{c}$ makes $O(T^{c}\operatorname*{len}(\operatorname*{conv}(S)))$ queries and satisfies
$\lim_{T\to\infty}\ \sup_{\mathbf{s}\in S^{T}}\ \sup_{\mu}\ \sup_{\pi^{*}:f(\pi^{*})\leq K}\ R_{T}(\mathbf{s},\mu,\pi^{*})=0$
Our algorithm does not need to know the Lipschitz constant $L$ or the number of segments; it only needs to know the time horizon $T$.
5.1 Intuition behind the algorithm
The algorithm maintains a set of buckets which partition the observed portion of the state space. Each bucket’s length determines the maximum loss in payoff we will allow from that subset of the state space. As long as the bucket contains a query from a prior time step, the Lipschitz constraint allows us to bound $\mu^{*}(s_{t})-\mu(s_{t},a_{t})$ based on the length of the bucket containing $s_{t}$. We always query if the bucket does not contain a prior query
The granularity of the buckets is controlled by a function $g$, with the initial buckets having length $1/g(T)$. Since we can expect one query per bucket, we need $g(T)\in o(T)$ to ensure sublinear queries.
Regardless of the bucket length, the adversary can still place multiple segments in the same bucket $B$. A single query only tells us the optimal action for one of those segments, so we risk a payoff as bad as $\mu^{*}(s_{t})-O(\operatorname*{len}(B))$ whenever we choose not to query. We can endure a limited number of such payoffs, but if we never query again in that bucket, we may suffer $\Theta(T)$ such payoffs. Letting $\mu^{*}(s_{t})=1$ for simplicity, that would lead to $\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\big{(}1-\frac{1}{O(g(T))}\big{)}^{\Theta(T)}$, which converges to 0 (i.e., guaranteed catastrophe) when $g(T)\in o(T)$.
This failure mode suggests a natural countermeasure: if we start to suffer significant (potential) losses in the same bucket, then we should probably query there again. One way to structure these supplementary queries is by splitting the bucket in half when enough time steps have involved that bucket. It turns out that splitting after $T/g(T)$ time steps is a sweet spot.
This approach is reminiscent of Algorithm 1 in that it collects an initial dataset (one query per initial bucket), and then performs supplemental queries as needed based on the observed states. We do not think this is a coincidence.
1:function AvoidCatastrophe($T\in\mathbb{N},\>g:\mathbb{N}\to\mathbb{N}$)
2:     $S_{Q}\leftarrow\emptyset$triangleright Previously queried states
3:     $\pi\leftarrow\emptyset$triangleright Records $\pi^{*}(s)$ for each $s\in S_{Q}$
4:     $\mathcal{B}\leftarrow\emptyset$triangleright The set of active buckets
5:     for $t$ from $1$ to $T$ do
6:         EvaluateState$(s_{t})$
7:function EvaluateState($s\in S$)
8:     if $s\not\in B$ for all $B\in\mathcal{B}$ then $\triangleright$ No bucket containing $s$: create a new bucket and try again
9:         $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $s\in B$
10:         $\mathcal{B}\leftarrow\mathcal{B}\cup\{B\}$
11:         $n_{B}\leftarrow 0$triangleright Number of time steps that have used $B$
12:         EvaluateState$(s)$
13:     else
14:         $B\leftarrow$ any bucket containing $s$
15:         if $S_{Q}\cap B=\emptyset$ then $\triangleright$ No queries in this bucket
16:              Query mentor and observe $\pi^{*}(s)$
17:              $\pi(s)\leftarrow\pi^{*}(s)$
18:              $S_{Q}\leftarrow S_{Q}\cup\{s\}$
19:              $n_{B}\leftarrow n_{B}+1$
20:         else if $n_{B}<T/g(T)$ then $\triangleright$ Bucket has a query and isn’t full: take that action
21:              Let $s^{\prime}\in S_{Q}\cap B$
22:              Take action $\pi(s^{\prime})$
23:              $n_{B}\leftarrow n_{B}+1$
24:         $\triangleright$ Bucket is full: split bucket and try again
25:              $B=[x,y]$
26:              $(B_{1},B_{2})\leftarrow\Big{(}\left[x,\frac{x+y}{2}\right],\left[\frac{x+y}{2},y\right]\Big{)}$
27:              $(s_{B_{1}},s_{B_{2}})\leftarrow(0,0)$
28:              $\mathcal{B}\leftarrow\mathcal{B}\cup\{B_{1},B_{2}\}\setminus B$
29:              EvaluateState$(s)$
Algorithm 2 achieves near-zero regret when the mentor’s policy has a bounded number of segments.
6 Limitations
Our paper contains both positive and negative results. As such, our claim in this paper is not “avoiding catastrophe is possible (or impossible) under assumptions X, Y, Z, and we argue that X, Y, Z hold”. Instead, we aim to understand the conditions under which it is possible to avoid catastrophe. Our claim is “avoiding catastrophe is possible under X, Y, Z, but impossible under A, B, C; we might sometimes expect X, Y, Z and sometimes expect A, B, C”. For example, some real-world instances might exhibit bounded Natarajan dimension, while others might not. Our paper elucidates what to expect from either scenario.
The goal, then, is to make the gap between X, Y, Z and A, B, C as small as possible: to isolate the boundary of tractability. Our results suggest that the key factor determining tractability is the complexity of the mentor policy class. Our negative result (Theorem 3.1) shows that a simple environment (bounded 1D state space, known uniform distribution over states, two actions) is not enough for tractability if the mentor policy class is complex. On the other hand, our primary positive result (Theorem 4.5) holds for quite general environments (unbounded $n$-dimensional state space, unknown arbitrary distribution over states, many actions), as long as the mentor policy class is simple.
One remaining gap between those two results is the assumption of $r$-locality in Theorem 4.5. Although we think $r$-locality is a fairly plausible assumption (as argued in Section 4), that assumption would need to be removed in order for Theorem 3.1 and Theorem 4.5 to fully match.
Another question is how far this positive result generalizes. In particular, the assumptions of (1) only caring about catastrophe and (2) i.i.d. states significantly limit the real-world applicability of our algorithm. In future work, we hope to extend this result to general MDPs.
Our secondary positive result (Theorem 5.2) also suggests that mentor policy complexity could be the key factor determining tractability. However, it has the drawback that the number of segments in the mentor policy is a less established measure of complexity than the well-studied notion of Natarajan dimension. In our opinion, the more serious limitation is the restriction to a 1D state space. Although Theorem 5.2 does permit adversarial state sequences, we feel that Theorem 4.5 does a better overall job of isolating the boundary of tractability.
7 Conclusion and future work
In this paper, we proposed a model of avoiding catastrophe in sequential decision-making. We showed that in general, any algorithm either constantly asks for help or is guaranteed to cause catastrophe. However, the problem becomes tractable when the mentor’s policy is relatively simple, for two different definitions of “simple”.
There remain some technical questions within this paper’s model: (1) Can Theorem 4.5 be generalized to adversarial state sequences? (2) Is $r$-locality necessary? (3) Are there are reasonable alternatives to the Lipschitz assumption? (Indeed, the utility of $r$-locality steps from the Lipschitz assumption.)
As mentioned, we are also excited to move beyond the current model to full MDPs: recall the open problem we posed in Section 1. Most applications are better captured by MDPs than by this paper’s assumptions of i.i.d. states and only caring about catastrophe.
Finally, although we intentionally avoided Bayesian inference in this paper, it seems premature to abandon that tool entirely. Are there tractable approximations to Bayesian approaches?
Acknowledgements
This work was supported in part by a gift from Open Philanthropy to the Center for Human-Compatible AI (CHAI) at UC Berkeley. This paper also benefited from discussions with many researchers at CHAI. We would like to particularly thank Michael Cohen, Bhaskar Mishra, Cam Allen, and Li Dayan for helpful feedback.
Appendix A Proof of Theorem 3.1
A.1 Proof roadmap
[…]
Recall that $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\mathbf{a}}$ denotes the expectation over both the random states and any additional randomness in the algorithm.
The proof proceeds via the following steps:
  1. 1.
    Prove that $f(T)=\sqrt{(|Q_{T}|+1)T}$ is asymptotically between $|Q_{T}|$ and $T$ (Lemma A.1).
  2. 2.
    Provide a simple variant of the Chernoff bound which we will apply multiple times (Lemma A.2).
  3. 3.
    Show that with high probability, $\sum_{j\in A}|V_{j}|$ is adequately large (Lemma A.3).
  4. 4.
    The key lemma is Lemma A.4, which shows that a randomly sampled $\mathbf{x}$ produces poor agent performance with high probability. The central idea is that at least $f(T)-|Q_{T}|$ sections are never queried (which is large, by Lemma A.1), so the agent has no way of knowing the optimal action in those sections. As a result, the agent picks the wrong answer at least half the time on average (and at least a quarter of the time with high probability). Lemma A.3 implies that a constant fraction of those time steps will have quite suboptimal payoffs, again with high probability.
  5. 5.
    To complete the proof, we observe that $\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu_{f,\mathbf{x}},\pi^{*})$. This is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\mathbf{x}}$ has high expected regret, then the worst case $\mu$ also has high expected regret.
Note that $\mathbf{s},\mathbf{a}$, and $\mathbf{x}$ are random variables, so all variables defined on top of them (such as $V_{j}$) are also random variables. In contrast, the partition $S=\{S_{1},\dots,S_{f(T)}\}$ and properties thereof (like the midpoints $m_{j}$) are not random variables.
A.2 Proof
Lemma A.1.
Let $a,b:\mathbb{N}\to\mathbb{N}$ be functions such that $a(x)\in o(b(x))$. Then $c(x)=\sqrt{a(x)b(x)}$ satisfies $a(x)\in o(c(x))$ and $c(x)\in o(b(x))$.
Proof.
Since $a$ and $b$ are strictly positive (and thus $c$ is as well), we have
$\frac{a(x)}{c(x)}=\frac{a(x)}{\sqrt{a(x)b(x)}}=\sqrt{\frac{a(x)}{b(x)}}=\frac{\sqrt{a(x)b(x)}}{b(x)}=\frac{c(x)}{b(x)}$
Then $a(x)\in o(b(x))$ implies
$\lim_{x\to\infty}\frac{a(x)}{c(x)}=\lim_{x\to\infty}\frac{c(x)}{b(x)}=\lim_{x\to\infty}\sqrt{\frac{a(x)}{b(x)}}=0$
as required. ∎
Lemma A.2.
Let $z_{1},\dots,z_{n}$ be i.i.d. variables in $\{0,1\}$ and let $Z=\sum_{i=1}^{n}z_{i}$. If $\operatorname*{\mathbb{E}}[Z]\geq M$, then $\Pr\big{[}Z\leq M/2\big{]}\leq\exp(-M/8)$.
Proof.
By the Chernoff bound for i.i.d. binary variables, we have $\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]\leq\exp(-\operatorname*{\mathbb{E}}[Z]/8)$. Since $-\operatorname*{\mathbb{E}}[Z]\leq-M$ and $\exp$ is an increasing function, we have $\exp(-\operatorname*{\mathbb{E}}[Z]/8)\leq\exp(-M/8)$. Also, $M/2\leq E[Z]/2$ implies $\Pr[Z\leq M/2]\leq\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]$. Combining these inequalities proves the lemma. ∎
Lemma A.3.
let $A\subseteq[f(T)]$ be any nonempty subset of sections. Then
$\Pr\left[\sum_{j\in A}|V_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T}{16f(T)}\right)$
Proof.
[…]
$\Pr\left[\sum_{j\in A}|V_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T|A|}{16f(T)}\right)\leq\exp\left(\frac{-T}{16f(T)}\right)$
with the last step due to $|A|\geq 1$. ∎
Lemma A.4.
Independently sample $\mathbf{x}\sim U(\{0,1\}^{f(T)})$ and $s\sim U(S^{T})$. [^13] That is, the entire set $\{x_{1},\dots,x_{f(T)},s_{1},\dots,s_{T}\}$ is mutually independent. Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$,
$\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq\exp\left(-\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}\right)$
Proof.
Part 1: setup. Let $J_{\neg Q}=\{j\in[f(T)]:s_{t}\not\in S_{j}\ \forall t\in Q_{T}\}$ be the set of sections that are never queried. Since each query appears in exactly one section (because each state appears in exactly one section), $|J_{\neg Q}|\geq f(T)-|Q_{T}|$.
For each $j\in J_{\neg Q}$, let $y_{j}$ be the action taken most frequently among time steps in $V_{j}$:
$y_{j}=\operatorname*{arg\,max}_{a\in\{0,1\}}\Big{|}\{t\in V_{j}:a=a_{t}\}\Big{|}$
Let $\bar{J}=\{j\in J_{\neg Q}:x_{j}\neq y_{j}\}$. For each $j\in\bar{J}$, let $V_{j}^{\prime}=\{t\in V_{j}:a_{t}\neq x_{j}\}$ be the set of time steps where the agent chooses the wrong action (assuming payoff function $\mu_{f,\mathbf{x}}$).
[…]
$\Pr\left[|\bar{J}|\leq\frac{f(T)-|Q_{T}|}{4}\right]\leq\exp\left(-\frac{f(T)-|Q_{T}|}{16}\right)$
Part 3: $V_{j}^{\prime}$ is not too small. Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in V_{j}$, so $a_{t}\in\{0,1\}$ for all $t\in V_{j}$. Since the agent chooses one of two actions for each $t\in V_{j}$, the more frequent action must be chosen chosen at least half of the time: $a_{t}=y_{j}$ for at least half of the time steps in $V_{j}$. Since $x_{j}\neq y_{j}$ for $j\in\bar{J}$, we have $a_{t}=y_{j}\neq x_{j}$ for those time steps, so $|V_{j}^{\prime}|\geq|V_{j}|/2$.
Part 4: a bound in terms of $\bar{J}$ and $V_{j}$. Consider any $j\in\bar{J}$ and $t\in V_{j}^{\prime}\subseteq V_{j}$. By definition of $V_{j}$, we have $|m_{j}-s_{t}|\leq\frac{1}{4f(T)}$. Then by definition of $\mu_{f,\mathbf{x}}$,
$\mu_{f,\mathbf{x}}(s_{t},a_{t})=$ $\ 1-L\left(\frac{1}{2f(T)}-|s_{t}-m_{j}|\right)$
$\leq$ $\ 1-L\left(\frac{1}{2f(T)}-\frac{1}{4f(T)}\right)$
$=$ $\ =1-\frac{L}{4f(T)}$
Now aggregating across time steps,
$\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ \prod_{j\in\bar{J}}\prod_{t\in V_{j}^{\prime}}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
$\leq$ $\ \prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}^{\prime}|}$
$\leq$ $\ \prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}|/2}$
[…]
$\log\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ \log\prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}|/2}$
$=$ $\ \sum_{j\in\bar{J}}\frac{|V_{j}|}{2}\log\left(1-\frac{L}{4f(T)}\right)$
$\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|V_{j}|}{8f(T)}$
[…]
$\log\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|V_{j}|}{8f(T)}$
$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T|\bar{J}|}{4f(T)}$
$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T}{4f(T)}\cdot\frac{f(T)-|Q_{T}|}{4}$
$=$ $\ -\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}$
Exponentiating both sides proves the lemma. ∎
Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$ for brevity. See 3.1
Proof.
If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{*}$. Let $f(T)=\sqrt{(g(T)+1)T}$. Then by Lemma A.1, $g(T)\in o(f(T))$ and $f(T)\in o(T)$. Combining this with $|Q_{T}|\leq g(T)$, we get $\lim_{T\to\infty}\alpha(T)=0$. Also, since $|Q_{T}|\in o(f(T))$, there exists $T_{0}$ such that $|Q_{T}|\leq f(T)/2$ for all $T\geq T_{0}$. Combining this with Lemma A.4 and noting that $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq 1$, we have
$\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ \>\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
$\leq$ $\ \alpha(T)\cdot 1+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}\right)$
$\leq$ $\ \alpha(T)+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LTf(T)/2}{2^{7}f(T)^{2}}\right)$
$=$ $\ \alpha(T)+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
whenever $T\geq T_{0}$. Since $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})=1$ always, we have [^14] Note that Fubini’s theorem means we need not worry about the order of the expectation operators.
$\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu_{f,\mathbf{x}},\pi^{*})$
$=\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})-\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\right]$
$\geq$ $\ 1-\alpha(T)-\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
Therefore
$\lim_{T\to\infty}\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$
$=$ $\ 1-0-(1-0)\cdot\exp(-\infty)$
[…]
as required. ∎
Appendix B Proof of Theorem 4.5
B.1 Proof roadmap
We use the following notation throughout the proof:
  1. 1.
    For each $t\in[T]$, let $X_{t},$ refer to the values of $X$ respectively at the start of time step $t$. Define $\pi_{t}$ similarly, except only for $t>g(T)$.
  2. 2.
    Let $V_{T}=\{t\in[T]:\mu(s_{t},a_{t})<\mu^{*}(s_{t})\}$ be the set of time steps with a suboptimal payoff.
  3. 3.
    Unless specified (e.g., $\Pr_{s\sim\mathcal{D}}$), all probabilities are with respect to $\mathbf{s}\sim\mathcal{D}^{T}$.
The proof proceeds in the following steps:
  1. 1.
    Show that $\pi_{t}$ meets the criteria for the fundamental theorem of multi-class PAC learning and thus has small expected error on a single state (Lemma B.1).
  2. 2.
    Use Lemma B.1 and a non-standard concentration inequality (Lemma B.2) to show that $|V_{T}|$ is small with high probability. This involves two steps (Lemmas B.3 and B.4).
  3. 3.
    Having a small number of mistakes is not sufficient, because a single payoff of $\mu(s_{t},a_{t})=0$ can ruin the entire endeavor. Lemmas B.5 and B.6 bound the suboptimality of a single payoff.
  4. 4.
    Lemma B.7 transforms the multiplicative objective into an additive form that is easier to manipulate.
  5. 5.
    Lemma B.8 bounds the total regret of the algorithm.
  6. 6.
    Lemma B.10 bounds the total number of queries based on a packing number lemma (Lemma B.9).
Lemmas B.8 and B.10 together imply Theorem 4.5.
B.2 Proof
Lemma B.1.
With probability at least $1-1/T$, we have $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ $\forall t>g(T)$.
Proof.
Let $\delta=1/T$. Then
$|X_{g(T)+1}|=$ $\ g(T)$
$=$ $\ \frac{cd}{\varepsilon(T)}\log(d\,T)$
$=$ $\ 2\frac{cd}{\varepsilon(T)}\log(d^{2}T^{2})$
$=$ $\ \frac{c}{\varepsilon(T)}(d\log(d^{2}T^{2})+d\log(d^{2}T^{2}))$
$\geq$ $\ \frac{c}{\varepsilon(T)}(d\log(d\,T\cdot T^{\frac{n}{n+0.5}})+\log T)$
$\geq$ $\ \frac{c}{\varepsilon(T)}\left(d\log(\frac{d|A|}{\varepsilon(T)})+\log\frac{1}{\delta}\right)$
Now consider any $t>g(T)$. By construction, $\pi_{t}$ is $(X_{t},r)$-local ($\pi$ and $X$ are not updated on every time step, but are always updated at the same time). Therefore $\pi_{t}$ is consistent with $X_{t}$ and in turn is consistent with $X_{g(T)+1}$ since $X_{g(T)+1}\subseteq X_{t}$. Since $X_{g(T)+1}$ is a set of i.i.d. samples of the requisite size, Theorem 4.4 implies that with at probability at least $1-1/T$, any policy $\pi\in\Pi$ consistent with $X_{g(T)+1}$ satisfies $\Pr_{s\sim\mathcal{D}}[\pi(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$. In particular, this holds for $\pi_{t}$ for all $t>g(T)$, completing the proof. ∎
Next, we wish to prove that $|V_{T}|$ is small. The $\{t\in V_{T}\}$ indicator variables are not independent, and standard martingale concentration bounds like the Azuma-Hoeffding inequality are not tight enough, so we use a specialized inequality from [8]. The statement below is slightly simplified from its original source, since we do not need the more general version.
Lemma B.2 (Corollary 2.1 in [8]).
Let $x_{0},\dots,x_{k}$ be a supermartingale and assume there exists $b\in(0,1]$ such that $x_{i}-x_{i-1}\in[-b,1]$ for all $i\in[k]$. Then for any $x>0$,
$\Pr[x_{k}\geq x]\leq\exp\left(-\frac{x^{2}}{2(kb+x/3)}\right)$
Lemma B.3.
Assume $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ for all $t>g(T)$. Then
$\Pr\left[|V_{T}|\geq 2T\varepsilon(T)\right]\leq\exp(-3T\varepsilon(T)/8)$
Proof.
Let $k=T-g(T)$ for brevity. For each $t\in\{g(T)+1,\dots,T\}$, define a random variable $y_{t-g(T)}=\mathbf{1}_{\pi_{t}(s_{t})\neq\pi^{*}(s_{t})}-\varepsilon(T)$. Also let $x_{0}=0$, and for $i\in[k]$, let $x_{i}=x_{i-1}+y_{i}$. This defines random variables $y_{1},\dots,y_{k}$ and $x_{0},\dots,x_{k}$. For each $i\in[k]$, let $y_{<i}=\{y_{1},\dots,y_{i-1}\}$ and $x_{<i}=\{x_{0},\dots,x_{i-1}\}$.
To show that $x_{0},\dots,x_{k}$ form a supermartingale, we first claim that $\operatorname*{\mathbb{E}}[y_{i}\mid y_{<i}]\leq 0$ for all $i\in[k]$. Regardless of what occurs on previous time steps, each state is sampled randomly from $\mathcal{D}$. Thus
$\operatorname*{\mathbb{E}}[y\mid y_{<i}]=\Pr_{s\sim\mathcal{D}}[\pi_{i+g(T)}\neq\pi^{*}(s)\mid y_{<i}]-\varepsilon(T)\leq 0$
Since $y_{<i}$ and $x_{<i}$ provide the same information (i.e., either can be constructed from the other), we have $\operatorname*{\mathbb{E}}[y_{i}\mid x_{<i}]\leq 0$. Therefore
$\operatorname*{\mathbb{E}}[x_{i}\mid x_{<i}]=$ $\ \operatorname*{\mathbb{E}}[x_{i-1}+y_{i}-\varepsilon(T)\mid x_{<i}]$
$=$ $\ \operatorname*{\mathbb{E}}[x_{i-1}\mid x_{<i}]+\operatorname*{\mathbb{E}}[y_{i}\mid x_{<i}]-\varepsilon(T)$
$\leq$ $\ x_{i-1}+\varepsilon(T)-\varepsilon(T)$
$=$ $\ x_{i-1}$
Also, since each $x_{i}$ is bounded, we have $\operatorname*{\mathbb{E}}[x_{i}]<\infty$ for all $i\in[k]$. Therefore $x_{0},\dots,x_{k}$ is indeed a supermartingale. Furthermore, $x_{i}-x_{i-1}=y_{i}\in[-\varepsilon(T),1]$ for all $i\in[k]$, so by Lemma B.2,
$\Pr[x_{k}\geq 2T\varepsilon(T)]\leq$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{2(k\varepsilon(T)+T\varepsilon(T)/3)}\right)$
Since $k\leq T$,
$\Pr[x_{k}\geq 2T\varepsilon(T)]\leq$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{2(T\varepsilon(T)+T\varepsilon(T)/3)}\right)$
$=$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{8T\varepsilon(T)/3}\right)$
$=$ $\ \exp(-3T\varepsilon(T)/8)$
as required. ∎
Lemma B.4.
We have $\Pr[|V_{T}|\geq 2T\varepsilon(T)]\leq\exp(-3T^{\frac{0.5}{n+0.5}}/8)+1/T$.
Proof.
Let $\xi$ be the event that $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ for all $t>g(T)$. Lemma B.1 states that $\Pr[\neg\xi]\leq 1/T$ and Lemma B.3 states that $\Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\xi]\leq\exp(-3T\varepsilon(T)/8)$. Then
$\Pr[|V_{T}|\geq 2T\varepsilon(T)]=$ $\ \Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\xi]\Pr[\xi]+\Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\neg\xi]\Pr[\neg\xi]$
$\leq$ $\ \exp(-3T\varepsilon(T)/8)\cdot 1+1\cdot 1/T$
$=$ $\ \exp(-3T^{1-\frac{n}{n+0.5}}/8)+1/T$
$=$ $\ \exp(-3T^{\frac{0.5}{n+0.5}}/8)+1/T$
as claimed. ∎
Remark B.1.
The astute reader note that Lemma B.5 is the only place in the paper where we invoke the assumption that the mentor’s policy is optimal. Both of our positive results (Theorems 4.5 and 5.2) can be generalized to a possibly suboptimal mentor policy $\pi^{m}$ if:
  1. 1.
    We modify the Lipschitz assumption to be $|\mu(s,\pi^{m}(s))-\mu(s,\pi^{m}(s^{\prime}))|\leq L||s-s^{\prime}||$. In words, the value of the mentor’s policy shouldn’t change too much between nearby states.
  2. 2.
    We define replace $\mu^{*}(s)$ with $\mu(s,\pi^{m}(s))$ throughout.
Lemma B.5 follows immediately from these two assumptions.
Also note that Lemmas B.5 and B.7 are not specific to this algorithm and will also be used in the proof of Theorem 5.2.
Lemma B.5.
If $||s-s^{\prime}||\leq\delta$ for some $\delta\geq 0$, then $\mu(s,\pi^{*}(s^{\prime}))\geq\mu^{*}(s)-2L\delta$.
Proof.
We have By the Lipschitz continuity of $\mu$, $\mu(s,\pi^{*}(s^{\prime}))\geq\mu(s^{\prime},\pi^{*}(s^{\prime}))-L||s-s^{\prime}||$. Also, $\mu(s^{\prime},\pi^{*}(s^{\prime}))=\max_{a\in A}\mu(s^{\prime},a)\geq\mu(s^{\prime},\pi^{*}(s))$ by assumption. Thus
$\mu(s,\pi^{*}(s^{\prime}))\geq$ $\ \mu(s^{\prime},\pi^{*}(s^{\prime}))-L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$)
$\geq$ $\ \mu(s^{\prime},\pi^{*}(s))-L||s-s^{\prime}||$ (By assumption, $\pi^{*}$ is optimal)
$\geq$ $\ \mu(s,\pi^{*}(s))-2L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$ again)
$=$ $\ \mu^{*}(s)-2L||s-s^{\prime}||$ (Definition of $\mu^{*}(s)$)
Plugging in $||s-s^{\prime}||\leq\delta$ completes the proof. ∎
Lemma B.6.
Under the conditions of Theorem 4.5, $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2Lr$ for all $t\in[T]$.
Proof.
If we query the mentor at time $t$, then $\mu(s_{t},a_{t})=\mu^{*}(s_{t})$, so assume we do not query the mentor at time $t$. Let $(s^{\prime},a^{\prime})=\operatorname*{arg\,min}_{(s,a)\in X_{t}}||s_{t}-s||$: then $||s_{t}-s^{\prime}||\leq r$. Since $\pi_{t}$ is $(X_{t},r)$-local by construction, there exists $(s^{\prime\prime},a^{\prime\prime})\in X_{t}$ such that $||s_{t}-s^{\prime\prime}||\leq r$ and $a_{t}=\pi_{t}(s_{t})=a^{\prime\prime}$. Then by Lemma B.5, $\mu(s_{t},\pi^{*}(s^{\prime\prime}))\geq\mu^{*}(s_{t})-2Lr$. Furthermore, because $(s^{\prime\prime},a^{\prime\prime})\in X$, we have $\pi^{*}(s^{\prime\prime})=a^{\prime\prime}=a_{t}$. Therefore $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2Lr$. ∎
Lemma B.7.
If $\mu^{*}(s_{t})\geq\mu_{0}>0$ and $c_{t}\in\big{[}0,\mu_{0}/2\big{]}$ for all $t\in[T]$, then
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq\prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
Proof.
Recall the standard limit $\lim_{\delta\to 0}\delta^{-1}\log(1+\delta)=1$. Thus for all $\varepsilon>0$, there exists $\delta_{0}>0$ such that whenever $|\delta|\leq\delta_{0}$, we have $|\delta^{-1}\log(1+\delta)-1|\leq\varepsilon$. For our purposes, $\varepsilon=0.5$ is sufficient, we only care about $\delta<0$, and we only need the lower bound. One can verify that $\delta_{0}=0.5$ is sufficient. To summarize, we have $1.5\delta\leq\log(1+\delta)$ whenever $|\delta|\leq 0.5$.
Let $\delta_{t}=-c_{t}/\mu^{*}(s_{t})$ for each $t\in[T]$. Then
$|\delta_{t}|=\frac{c_{t}}{\mu^{*}(s_{t})}\leq\frac{\mu_{0}/2}{\mu_{0}}=0.5$
and so
$-\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\left(1-\frac{c_{t}}{\mu^{*}(s_{t})}\right)$
$\log(\mu^{*}(s_{t}))-\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\left(\mu^{*}(s_{t})-c_{t}\right)$
Now summing over $t$,
$\sum_{t=1}^{T}\log(\mu^{*}(s_{t}))-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \sum_{t=1}^{T}\log\left(\mu^{*}(s_{t})-c_{t}\right)$
$\log\prod_{t=1}^{T}\mu^{*}(s_{t})-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
Exponentiating both sides yields
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq\exp\left(-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\right)\prod_{t=1}^{T}\mu^{*}(s_{t})$
Applying the standard inequality $e^{x}\geq 1+x$ gives us
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq$ $\ \left(1-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\right)\prod_{t=1}^{T}\mu^{*}(s_{t})$
$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}\frac{\prod_{i=1}^{T}\mu^{*}(s_{i})}{\mu^{*}(s_{t})}$
Since $\mu^{*}(s_{t})\in[0,1]$ for all $t\in[T]$, $\prod_{i=1}^{T}\mu^{*}(s_{i})\leq\mu^{*}(s_{t})$. Thus
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}\frac{\mu^{*}(s_{t})}{\mu^{*}(s_{t})}$
$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
as required. ∎
Lemma B.8.
Under the conditions of Theorem 4.5, Algorithm 1 satisfies
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq\frac{6L}{T^{\frac{0.5}{n+0.5}}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
Proof.
Let $\mu_{0}=4Lr$. If there exists $t\in[T]$ with $\mu^{*}(s_{t})<\mu_{0}$, then
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})<\mu_{0}-0=4L\varepsilon(T)^{1/n}=4LT^{\frac{-1}{n+0.5}}\leq 6LT^{\frac{-0.5}{n+0.5}}$
so the claim is satisfied. Thus assume $\mu^{*}(s_{t})\geq\mu_{0}$ for all $t\in[T]$. Let $c_{t}=0$ for $t\not\in V_{T}$ and $c_{t}=2Lr$ for $t\in V_{T}$. Then by Lemma B.6,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})=$ $\ \prod_{t\in V_{T}}\mu(s_{t},a_{t})\prod_{t\not\in V_{T}}\mu(s_{t},a_{t})$
$\geq$ $\ \prod_{t\in V_{T}}\left(\mu^{*}(s_{t})-2Lr\right)\prod_{t\not\in V_{T}}\mu^{*}(s_{t})$
$=$ $\ \prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
Since $c_{t}\in[0,\mu_{0}/2]$ for all $t\in[T]$, Lemma B.7 implies that
$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-3L|V_{T}|r$
If $|V_{T}|<2T\varepsilon(T)$, then
$R_{T}(\mathbf{s},\mu,\pi^{*})=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
$\leq$ $\ 6LT\varepsilon(T)r$
$=$ $\ 6LT\varepsilon(T)^{(n+1)/n}$
$=$ $\ 6LT^{(1+\frac{-n}{n+0.5}\cdot\frac{n+1}{n})}$
$=$ $\ 6LT^{(\frac{n+0.5}{n+0.5}-\frac{n+1}{n+0.5})}$
$=$ $\ 6LT^{\frac{-0.5}{n+0.5}}$
Then with the help of Lemma B.4 (and noting that $R_{T}(\mathbf{s},\mu,\pi^{*})\leq 1$ always),
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq$ $\ \Pr[|V_{T}|\geq 2T\varepsilon(T)]\cdot 6LT^{\frac{-0.5}{n+0.5}}+\Pr[|V_{T}|\geq 2T\varepsilon(T)]\cdot 1$
$\leq$ $\ 6LT^{\frac{-0.5}{n+0.5}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
Definition B.2.
Let $(V,||\cdot||)$ be a normed vector space and let $\delta>0$. Then $W\subseteq V$ is a $\delta$-packing of $V$ if for all $x,y\in W$, $||x-y||>\delta$. The $\delta$-packing number of $V$, denoted $\mathcal{M}(V,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $V$.
Lemma B.9 (Theorem 14.2 in [32]).
Let $B$ be the unit ball and let |||||$\cdot$||| | ⋅ | | denote Euclidean distance. Assume $V\subset\mathbb{R}^{n}$ is convex, and $\delta B\subset V$ for some $\delta>0$. Then
$\mathcal{M}(V,||\cdot||,\delta)\leq\left(\frac{3}{\delta}\right)^{n}\frac{\operatorname*{vol}(V)}{\operatorname*{vol}(B)}$
Lemma B.10.
[…]
Proof.
Let $\tilde{S}=\{s_{t}:t\in Q_{T}\text{ and }t>g(T)\}$ be the set of states queried after $t=g(T)$. Since we always query on the first $g(T)$ time steps, the total number of queries is $g(T)+|\tilde{S}|$.
We claim that for any $s,s^{\prime}\in\tilde{S}$, we have $||s-s^{\prime}||>r$. Suppose not, and WLOG assume $s$ was queried after $s^{\prime}$, with $s=s_{t}$. Since $s^{\prime}$ was queried earlier, $(s^{\prime},\pi^{*}(s^{\prime}))\in X_{t}$. Therefore
$\min_{(s^{\prime\prime},a)\in X_{t}}||s-s^{\prime\prime}||\leq||s-s^{\prime}||\leq r$
But then we would have taken the action $\pi_{t}(s_{t})$ instead of querying, which is a contradiction. Hence $\tilde{S}$ is a $r$-packing of $\operatorname*{conv}(\mathbf{s})$, and thus is also an $r$-packing of $\operatorname*{conv}(\mathbf{s}\cup rB)$. Since $rB\subset\operatorname*{conv}(\mathbf{s}\cup rB)$, Lemma B.9 implies that
$|\tilde{S}|\leq\mathcal{M}(\operatorname*{conv}(\mathbf{s}\cup rB),||\cdot||,r)\leq\left(\frac{3}{r}\right)^{n}\frac{\operatorname*{vol}(\operatorname*{conv}(\mathbf{s}\cup rB))}{\operatorname*{vol}(B)}=\frac{3^{n}\operatorname*{vol}(\operatorname*{conv}(\mathbf{s}\cup rB))}{\varepsilon(T)\operatorname*{vol}(B)}$
Theorem 4.5 follows from Lemmas B.8 and B.10.
Appendix C Proof of Theorem 5.2
C.1 Notation for the proof
We will use the following notation throughout the proof of Theorem 5.2:
  • Let $V_{T}=\{t\in[T]:\mu(s_{t},a_{t})<\mu^{*}(s)\}$ be the set of time steps with a suboptimal payoff.
  • Let $B_{t}$ be the bucket that is used on time step $t$ (as defined on line 14 of Algorithm 2).
  • Let $d(B)$ be the depth of bucket $B$
    • Buckets created on line 9 are depth 0.
    • We refer to $B_{1},B_{2}$ created on line 26 as the children of the bucket $B$ defined on line 14.
    • If $B^{\prime}$ is the child of $B$, $d(B^{\prime})=d(B)+1$.
[…]
  • Viewing the set of buckets are a binary tree defined by the “child” relation, we use the terms “ancestor” and ”descendant” in accordance with their standard tree definitions.
  • Let $\mathcal{B}_{V}=\{B:\exists t\in V_{T}\text{ s.t. }B_{t}=B\}$ be the set of buckets that ever produced a suboptimal payoff.
  • Let $\mathcal{B}_{V}^{\prime}=\{B\in\mathcal{B}_{V}:\text{no descendant of $B$ is in $\mathcal{B}_{V}$}\}$.
C.2 Proof roadmap
The proof proceeds in the following steps:
  1. 1.
    Bound the total number of buckets and therefore the total number of queries (Lemma C.1).
  2. 2.
    Bound the suboptimality on a single time step based on the bucket length and the Lipschitz constant (Lemma C.2).
  3. 3.
    Bound the sum of bucket lengths on time steps where we make a mistake (Lemma C.4), with Lemma C.3 as an intermediate step. This captures the total amount of suboptimality.
  4. 4.
    The bound from Lemma C.4 is in additive form, whereas the objective is multiplicative. Lemma B.7 factors the objective into one term for the optimal payoffs and one term for the suboptimality in the form of Lemma C.4.
  5. 5.
    Combine Lemmas C.2, C.4 and B.7 to bound the total regret (Lemma C.5).
  6. 6.
    Prove Theorem 5.2 using Lemma C.1, Lemma C.5, and some light arithmetic.
C.3 Proof
Lemma C.1.
Algorithm 2 performs at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries.
Proof.
Algorithm 2 performs at most one query per bucket, so the total number of queries is bounded by the total number of buckets. There are two ways to create a bucket: from scratch (line 9), or by splitting an existing bucket (line 26).
Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|s_{t}-s_{t^{\prime}}|=g(T)\operatorname*{len}(\operatorname*{conv}(\mathbf{s}))$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}s_{t},\max_{t\in[T]}s_{t}]$. At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{len}(\operatorname*{conv}(\mathbf{s}))+2$.
We split a bucket $B$ when $n_{B}$ reaches $T/g(T)$, which creates two new buckets. Since each time step increments $n_{B}$ for a single bucket $B$, and there are a total of $T$ time steps, the total number of buckets created via splitting is at most
$\frac{2T}{T/g(T)}=2g(T)$
Therefore the total number of buckets ever in existence is $(2+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)+2\leq(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$, so Algorithm 2 performs at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries. ∎
Lemma C.2.
For each $t\in[T]$, $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2L\operatorname*{len}(B_{t})$.
Proof.
If we query the mentor at time $t$, $\mu(s_{t},a_{t})=\mu^{*}(s_{t})$. Thus assume we do not query the mentor at time $t$: then there exists $s^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 2) such that $a_{t}=\pi(s^{\prime})=\pi^{*}(s^{\prime})$. Since $s_{t}$ and $s^{\prime}$ are both in $B_{t}$, $|s_{t}-s^{\prime}|\leq\operatorname*{len}(B_{t})$. Applying Lemma B.5 completes the proof. ∎
Lemma C.3.
If $\pi^{*}$ has at most $K$ segments, $|\mathcal{B}_{V}^{\prime}|\leq K$.
Proof.
If two buckets $B$ and $B^{\prime}$ overlap by more than at a single point, one must be the descendant of the other. Thus by definition of $\mathcal{B}_{V}^{\prime}$, any two buckets in $\mathcal{B}_{V}^{\prime}$ overlap by at most a single point.
Now consider any $B\in\mathcal{B}_{V}^{\prime}$. By definition, there exists $t\in V_{T}$ such that $s_{t}\in B$. Then there exists $s^{\prime}\in B$ (as defined in Algorithm 2) such that $a_{t}=\pi(s^{\prime})=\pi^{*}(s^{\prime})$. Since $t\in V_{T}$, we have $\pi^{*}(s_{t})\neq a_{t}=\pi^{*}(s^{\prime})$. Thus $s_{t}$ and $s^{\prime}$ are in different segments; let $S_{B}$ be the segment containing $\min(s_{t},s^{\prime})$. Then $\min(s_{t},s^{\prime})\leq\max(S_{B})<\max(s_{t},s^{\prime})\leq\max(B)$.
Consider any $B,B^{\prime}\in\mathcal{B}_{V}^{\prime}$ with $B\neq B^{\prime}$ and WLOG assume $\max(B)<\max(B^{\prime})$. Then $\max(S_{B})<\max(B)\leq\max(S_{B}^{\prime})$, so $S_{B}\neq S_{B^{\prime}}$. Thus the number of segments in $\pi^{*}$ is at least $\mathcal{B}_{V}^{\prime}$. Since the number of segments in $\pi^{*}$ is at most $K$, we must have $|\mathcal{B}_{V}^{\prime}|\leq K$. ∎
Lemma C.4.
[…]
Proof.
For every $t\in V_{T}$, we have $B_{t}=B$ for some $B\in\mathcal{B}_{V}$, so
$\sum_{t\in V_{T}}\operatorname*{len}(B_{t})=\sum_{B\in\mathcal{B}_{V}}\ \sum_{t\in V_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
Next, observe that every $B\in\mathcal{B}_{V}\setminus\mathcal{B}_{V}^{\prime}$ must have a descendent in $\mathcal{B}_{V}^{\prime}$: otherwise we would have $B\in\mathcal{B}_{V}^{\prime}$. Let $\mathcal{A}(B)$ denote the set of ancestors of $B$, plus $B$ itself. Then we can write
$\sum_{t\in V_{T}}$ $\ \operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ \sum_{t\in V_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
$=$ $\ \sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ |\{t\in V_{T}:B=B_{t}\}|\cdot\operatorname*{len}(B_{t})$
[…]
$\sum_{B\in\mathcal{A}(B^{\prime})}$ |$\ \frac{|\{t\in V_{T}:B=B_{t}\}|}{g(T)2^{d(B)}}\leq\frac{T}{g(T)^{2}}\sum_{B\in\mathcal{A}(B^{\prime})}\ \frac{1}{2^{d(B)}}$
$=$ $\ \frac{T}{g(T)^{2}}\sum_{d=0}^{d(B^{\prime})}\frac{1}{2^{d}}\leq\frac{T}{g(T)^{2}}\sum_{d=0}^{\infty}\frac{1}{2^{d}}=\frac{2T}{g(T)^{2}}$
Then by Lemma C.3,
$\sum_{t\in V_{T}}\operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\frac{2T}{g(T)^{2}}=\frac{2T|\mathcal{B}_{V}^{\prime}|}{g(T)^{2}}\leq\frac{2KT}{g(T)^{2}}$
as claimed. ∎
Lemma C.5.
For any $\mathbf{s}\in S^{T}$, any Lipschitz $\mu$, any $\pi^{*}$ with at most $K$ segments, and any $g:\mathbb{N}\to\mathbb{N}$, Algorithm 2 satisfies
$R_{T}(\mathbf{s},\mu,\pi^{*})\leq\frac{6L}{g(T)}\max\left(1,\frac{TK}{g(T)}\right)$
Proof.
Let $\mu_{0}=4L/g(T)$. Suppose there exists $t\in[T]$ such that $\mu^{*}(s_{t})<\mu_{0}$. Then
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\mu_{0}-0\leq\frac{6L}{g(T)}$
so the claim is satisfied. Thus assume $\mu^{*}(s_{t})\geq\mu_{0}$ for all $t\in[T]$. Let $c_{t}=0$ for $t\not\in V_{T}$ and $c_{t}=2L\operatorname*{len}(B_{t})$ for $t\in V_{T}$. Then by Lemma C.2,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})=$ $\ \prod_{t\in V_{T}}\mu(s_{t},a_{t})\prod_{t\not\in V_{T}}\mu(s_{t},a_{t})$
$\geq$ $\ \prod_{t\in V_{T}}\left(\mu^{*}(s_{t})-2L\operatorname*{len}(B_{t})\right)\prod_{t\not\in V_{T}}\mu^{*}(s_{t})$
$=$ $\ \prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
Note that $\operatorname*{len}(B_{t})\leq 1/g(T)$ for all $t\in[T]$, so $c_{t}\in[0,\mu_{0}/2]$. Thus by Lemma B.7,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu(s_{t},a_{t})-1.5\sum_{t=1}^{T}c_{t}$
$=$ $\ \prod_{t=1}^{T}\mu(s_{t},a_{t})-3L\sum_{t\in V_{T}}\operatorname*{len}(B_{t})$
Then by Lemma C.4,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-\frac{6LKT}{g(T)^{2}}$
and so
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\frac{6LKT}{g(T)^{2}}$
as required. ∎
See 5.2
Proof.
Lemma C.1 implies the query bound. Since $KT\geq T\geq g(T)$, Lemma C.5 implies that for any such $\mathbf{s},\mu$, and $\pi^{*}$,
$R_{T}(\mathbf{s},\mu,\pi^{*})\leq$ $\ \frac{6L}{g(T)}\max\left(1,\frac{TK}{g(T)}\right)=\frac{6L}{g(T)}\cdot\frac{TK}{g(T)}$
$=$ $\ \frac{6LTK}{T^{2c}}=6LKT^{1-2c}$